OpenStudy (anonymous):

Bikesville is 200 miles from Restsville. Jasmine can complete the entire trip in 8 hours, and Monte takes two hours longer than Jasmine to complete the entire trip. If one begins in Bikesville and the other in Restville, then approximately how long will it take them to meet?

OpenStudy (anonymous):

4 hours

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Jamie can finish the trip in 8 hrs: 200/8 = 25 mph. Monte takes 2 hrs longer, or 10 hrs: 200/10 = 20. their combined rates, are pedaling towards each other: 200/45 = 4 4/9 , which is a little less than 4.5. The question is asking for an approximation, so that gives us: (C) 4hrs

OpenStudy (anonymous):

where u get the 45

OpenStudy (anonymous):

20+25=45